# Out-of-Context Misinformation Detection via Variational Domain-Invariant Learning with Test-Time Training

### v1

@@ Abstract
Out-of-context misinformation (OOC) is a low-cost form of misinformation in news reports, which refers to place authentic images into out-of-context or fabricated image-text pairings. This problem has attracted significant attention from researchers in recent years. Current methods focus on assessing image-text consistency or generating explanations. However, these approaches assume that the training and test data are drawn from the same distribution. When encountering novel news domains, models tend to perform poorly due to the lack of prior knowledge. To address this challenge, we propose VDT to enhance the domain adaptation capability for OOC misinformation detection by learning domain-invariant features and test-time training mechanisms. Domain-Invariant Variational Align module is employed to jointly encodes source and target domain data to learn a separable distributional space domain-invariant features. For preserving semantic integrity, we utilize domain consistency constraint module to reconstruct the source and target domain latent distribution. During testing phase, we adopt the test-time training strategy and confidence-variance filtering module to dynamically updating the VAE encoder and classifier, facilitating the model’s adaptation to the target domain distribution Extensive experiments conducted on the benchmark dataset NewsCLIPpings demonstrate that our method outperforms state-of-the-art baselines under most domain adaptation settings.

### v2

@@ Abstract
Out-of-context misinformation (OOC) is a low-cost form of misinformation in news reports, which refers to place authentic images into out-of-context or fabricated image-text pairings. This problem has attracted significant attention from researchers in recent years. Current methods focus on assessing image-text consistency or generating explanations. However, these approaches assume that the training and test data are drawn from the same distribution. When encountering novel news domains, models tend to perform poorly due to the lack of prior knowledge. To address this challenge, we propose VDT to enhance the domain adaptation capability for OOC misinformation detection by learning domain-invariant features and test-time training mechanisms. Domain-Invariant Variational Align module is employed to jointly encodes source and target domain data to learn a separable distributional space domain-invariant features. For preserving semantic integrity, we utilize domain consistency constraint module to reconstruct the source and target domain latent distribution. During testing phase, we adopt the test-time training strategy and confidence-variance filtering module to dynamically updating the VAE encoder and classifier, facilitating the model’s adaptation to the target domain distribution. Extensive experiments conducted on the benchmark dataset NewsCLIPpings demonstrate that our method outperforms state-of-the-art baselines under most domain adaptation settings.
@@ Conclusion
Appendix D Appendix D. Sensitivity analysis of CVF
To evaluate the robustness of the proposed confidence-variance filtering mechanism in the test-time training stage, we conduct a sensitivity analysis on its two key parameters $\alpha_{1},\alpha_{2}$. Specifically, we vary the scaling factors of confidence and the variance to observe their influence on model adaptation performance. The experimental results are shown in Table 5.
$\alpha_{1},\alpha_{2}$ $\rightarrowB U,W\rightarrowG$ $\rightarrowU B,G\rightarrowW$
F1F1/Acc F1F1/Acc F1F1/Acc F1F1/Acc
2, -1 72.26/72.30 76.59/76.63 81.96/81.95 79.02/79.02
1, 1 72.09/72.40 76.43/76.49 80.88/80.89 78.73/78.73
1, 0 71.95/72.07 76.65/76.68 81.72/81.73 78.86/78.90
0, 1 71.80/72.23 76.60/76.65 80.89/80.90 78.66/78.67
w/o gate 71.28/71.36 76.05/76.05 80.75/80.78 78.18/78.23
Table 5: Performance (%) of different settings of $\alpha s$ on NewsCLIPpings. In X $\rightarrow$ Y, X denotes the source domain, Y denotes the target domain.
The results demonstrate that the original setting (2, –1) consistently achieves the best or near-best performance across all four domain settings. This configuration emphasises high confidence (positive weight) while penalizing high variance (negative weight). The strong performance indicates that samples with confident and stable predictions indeed provide more reliable gradients during test-time training stage. When both confidence and variance are positively weighted (1, 1), the gate mechanism may incorrectly assign high importance to uncertain samples (high variance), weakening the adaptation signal. This explains the slight performance drop. Both single-factor filtering (1, 0) and (0, 1) maintain reasonable performance, but neither outperforms the combined setting. This demonstrates that confidence and variance capture complementary aspects of sample reliability, and omitting either one reduces adaptation effectiveness. However, removing the gate leads to the most obvious degradation.
In summary, the experimental results show that combining high confidence and low variance produces the most effective sample selection strategy for test-time training, while the mechanism exhibits good robustness to parameter variations.
Appendix E Appendix E. MMD Distance Comparison
To evaluate the effectiveness of the proposed DIVA module in reducing domain discrepancy, we measure the Maximum Mean Discrepancy (MMD) between the source-domain and target-domain feature distributions. Specifically, we compute the MMD values before and after applying the DIVA module during model training. Four domain-transfer pairs ($\rightarrowB, UW\rightarrowG$, $\rightarrowU, BG\rightarrowW$) from the NewsCLIPpings dataset are included for evaluation. This experiment aims to verify whether the proposed DIVA module in VDT can effectively align the feature spaces of source and target domains, thereby facilitating better domain-invariant representation learning.
MMD $\rightarrowB UW\rightarrowG$ $\rightarrowU BG\rightarrowW$
before 0.0607 0.0464 0.0529 0.0463
after 1.72e-4 1.18e-4 1.97e-6 3.68e-4
Table 6: The MMD distance between source-domain and target-domain feature distributions. In X $\rightarrow$ Y, X denotes the source domain, Y denotes the target domain.
Table 6 presents the MMD values before and after applying the DIVA module. It can be known from the experimental results that MMD is dramatically reduced across all domain pairs. The domain-invariant variational align module is designed to enforce shared semantic structure between source and target domains, the empirical reduction in MMD directly supports this theoretical claim.
Appendix F Appendix F. Evaluating Dependency on Source Domain
To investigate whether our model relies excessively on any specific source domain, we conduct a source domain ablation study. Under the original multi-source setting (use all domains except the target domain as the source domain), the model is trained normally. Then, we removes one source domain while keeping all other settings identical. For each configuration, we evaluate the final performance on all target domain setups, reporting $F1$ score in figure 6.
Figure 6: The $F1$ score comparison of the model after removing the source domain in sequence. In $\rightarrowY, X$ denotes the source domains, Y denotes the target domain. And ”w/o B” means removes ”B” (i.e. bbc) domain data.
[…]
Overall, the consistently small performance variations across all ablation settings indicate that the model does not overfit to or rely disproportionately on any single source domain. Instead, it effectively leverages information from all available sources while maintaining generalization capability when one domain is absent.